# Analytic solutions for the Bianchi I universe coupled to several barotropic perfect fluids

### v2

@@ 4 Explicit particular solutions
  • For cases with $\{0<\rho_{0+}$ and $0<\rho_{0-}\}$ or $\{\rho_{0+}<0$ and $\rho_{\rm threshold}<\rho_{0-}\}$, the potential has no positive roots, and the evolution of the scale factor is given by (4.13) or (4.14), respectively. These solutions describe an indefinite expansion $a\in(0,+\infty)$, with the domain $\tau\in(-\infty,\tau_{0})$, where
    $\tau_{0}:=-\frac{1}{3P(1-\omega_{+})}\ln\left({\rm sgn}(\rho_{0+})+\frac{2\sqrt{3}P\sqrt{\rho_{0-}}}{|\rho_{0+}|\sqrt{\kappa}}\right),$
    is the time when infinite volume is reached, while the singularity is located at $\tau\to-\infty$. A particular case of this solution corresponds to the case analyzed in Ref. [3], with dust ($\omega_{+}=0$ and $\rho_{0+}=\rho_{0{\rm dust}}$), a positive cosmological constant ($\omega_{-}=-1$ and $\rho_{0-}=\Lambda/\kappa$), and stiff matter (included as $P\to(P^{2}+\frac{\kappa}{3}\rho_{0{\rm stiff}})^{1/2}$).
  • For cases with $\rho_{0-}<0$ and $P\neq 0$, the potential $V(a)$ has a single positive root, the solution is given either by (4.13) if $\rho_{0+}>0$, or by (4.14) if $\rho_{0+}<0$, with domain $\tau\in\mathbb{R}$. The range of $a$ is $a\in(0,a_{\max}]$, and thus there is a recollapse at
@@
  • For cases with $\{\rho_{0+}<0$ and $0\leq\rho_{0-}\leq\rho_{\rm threshold}\}$ the potential $V(a)$ has two positive roots,
@@
    On the one hand, a finite universe described by the functional form (4.14), with image $a\in(0,a_{-}]$, which thus presents an initial and a final singularity, as well as a recollapse at $a_{-}$. On the other hand, a singularity-free universe, with an evolution described by the functional form (4.15), which leads to an image $a\in[a_{+},+\infty)$, and thus undergoes a bounce at $a=a_{+}$. An interesting property of this cosmology is that, even if one needs to assume an exotic fluid with $\rho_{0+}<0$ to resolve the singularity, for large volumes $a\to+\infty$ the density of such fluid decays until becoming negligible, and the other nonexotic component $\rho_{-}>0$ dominates completely the evolution. Let us also mention that, in the isotropic case $P\to 0$, the threshold energy $\rho_{\rm threshold}$ tends to infinity and thus, the exotic $\rho_{0+}$ species is able to produce a bounce and provide a singularity-free cosmological evolution, independently of the energy contained in the $\rho_{0-}$ component.
    In the degenerate case with $\rho_{0-}=\rho_{\rm threshold}$, both roots coincide $a_{+}=a_{-}$, $V(a)$ presents a maximum there, and thus $a=a_{+}=a_{-}$ is an unstable equilibrium point. This implies that, unlike in the general case that the system reaches $a_{\pm}$ in finite time, in the degenerate case $a_{\pm}$ will be reached only in an infinite amount of time.
@@ 5 Asymptotic behavior
Let us now study the asymptotic behavior, both towards the singularity and towards large volumes, of the Bianchi I model filled with several barotropic fluids. As commented in Subsec. 2.2, due to their different barotropic indices, the scaling of the species differ. For $a\to 0$, the fluid with the maximum barotropic index $\omega_{+}$ will dominate, while for $a\to+\infty$ all the species will be negligible except the one with the minimum $\omega_{-}$. Therefore, we will consider the regime where only one fluid is dominant, so we can then use the explicit solution (4.9). Since recollapsing or bouncing cosmologies with exotic matter fields may not reach the limit of either large or small volumes, for simplicity, here we will assume that in both limits the corresponding dominant species has a positive energy density $\rho_{0}$.
As commented in the previous section, we will analyze the curvature by means of the Kretschmann scalar (2.13) and the Weyl square (2.14), which are related by equation (2.18). Specifically, considering a universe with a single species with $\rho_{0}>0$ and $\omega\in[-1,1)$, which follows the evolution (4.9), for $P\neq 0$ the Kretschmann scalar reads
@@ 5.1 Isotropization towards large volumes
where $a_{0}:=\left[\frac{3}{4}\kappa\rho_{0}(1-\omega)^{2}\right]^{-\frac{1}{3(1-\omega)}}$, and one should take into account that the evolution (4.9) is only valid for $\omega\in[-1,1)$ and $\rho_{0}>0$, and thus also this approximation. Moreover, taking this limit in the evolution (3.6)–(3.7) of the shape parameters leads to
@@
5.2 Asymptotics towards the singularity and blow-up of the curvature
@@
for $P\neq 0$ and $\omega\neq 1$. For the isotropic ($P=0$) case with $\omega\neq 1$, the evolution is given by (4.10). Hence, unless $P=0$, in this limit $a$ evolves exponentially in $\tau$, mirroring precisely the evolution of the vacuum (Kasner) solution, as given in Eq. (4.2). In order to translate this to in terms of the cosmological time $t$, we can just consider the relation obtained for the vacuum solution (4.3), which leads to the form
@@
To finish with this analysis, let us evaluate the Kretschmann scalar (5.1) in this limit. For the $P=0$ case, it scales as a power-law (5.2), and therefore always diverges towards the singularity, regardless of the matter type. For $\rho_{0}>0$, $\omega\neq 1$, and $P\neq 0$ in the limit towards the singularity, $\tau\to-\infty$,
@@
As we can observe from (5.18), in terms of the cosmological time, the form of the Kretschmann scalar in this limit does not depend on the matter content. In fact, its scaling follows the same power-law as in the large-volume limit (5.12). But, since the singularity is located at $t=0$, in this case the scalar will diverge.
Note that in the above expressions (5.17)–(5.19), we have only displayed the leading term, which, if $\sin(3\theta/2)=0$, will be vanishing. However, under the presence of matter, there are always subdominant diverging terms, and the Kretschmann scalar generically diverges. The vacuum scenario is the only case for which there are certain trajectories that have a constant and exactly vanishing Kretschmann scalar, which thus does not diverge towards the singularity. It is easy to see this feature from its exact form (5.3): if either $p_{+}=-P$ (and thus $\theta=0$) or $p_{+}=P/2$ (and thus $\theta=2\pi/3$ or $\theta=4\pi/3$), then $K=0$. These special trajectories correspond to the following three sets of Kasner exponents (4.6): $(p_{1}=p_{2}=0,p_{3}=1)$, $(p_{1}=p_{2}=0,p_{2}=1)$, and $(p_{2}=p_{3}=0,p_{1}=1)$, which define the only three possibilities with all the Kasner exponents being nonnegative, and it implies that none of the scale factors (4.5) expands towards the singularity.
@@ 6 Conclusions
We have obtained exact analytic solutions to describe the evolution of the Bianchi I geometry coupled to several perfect fluid species. The dynamical variables that describe the evolution of the model are the energy density of each species $\rho_{s}$, the two shape parameters $\beta_{\pm}$, that encode the anisotropy, and the average scale factor $a$, which is defined as the geometric average of the scale factors in the different spatial directions.
@@
In the chosen gauge, the evolution of the matter energy density $\rho_{s}$ and the shape parameters $\beta_{\pm}$ can be analytically obtained in general. Furthermore, the average scale factor $a$ follows the Friedmann equation (3.8), where the anisotropies are encoded in a constant term $P^{2}=(p_{+}^{2}+p_{-}^{2})$, which can be understood as the (square of the) momentum of a massless scalar field. The solution to this equation can be reduced to a quadrature, and we have described the qualitative evolution of the system by understanding the Friedmann equation as an energy equation for a particle moving on a potential [see Eq. (3.11)]. In addition, we have obtained the explicit analytic solution for a number of cases. In particular, we have derived the Kasner solution for vacuum (4.2), as well as the solution for a single fluid species (4.9). Although this solution has been studied in the literature, we have not found it anywhere given in such explicit and compact form as here. We have also considered the case with a negative energy density (4.11), which includes the model with a negative cosmological constant, and generically produces a recollapse of the universe, leading to a cosmology with an initial and a final singularity. Concerning the case with two fluid species, the solution is obviously more complicated, but, if the barotropic indices obey certain relation, the integral can be performed and the solution can be explicitly written. Such relation includes the particular cases of a Bianchi I universe either with cosmological constant and dust, or with radiation and cosmic strings. For completeness, we have obtained the explicit solution (4.13)–(4.15) for all the different signs of the energy densities. Exotic species with a negative sign of the density can produce turning points of the scale factor and bound (either from above or below) its possible values. Among the different possibilities, we find an interesting case, with an exotic and nonexotic component, that leads to a singularity-free cosmology: the universe collapses from infinite volume until certain minimum value of the scale factor, where a bounce happens, and then it expands forever. The exotic component is dominant around the bounce, while, for large volumes, it decays and becomes negligible as compared to the nonexotic one. We consider that this, and similar models, could be worth studying as an effective description of an nonsingular cosmology.
@@
Finally, we have analyzed the asymptotic behavior of the Bianchi I universe for both large volumes and towards the singularity. In such limits, we have studied in detail the scaling of the different curvature invariants, and explicitly derive their dependence on the different parameters of the model. For large volumes, the anisotropies decay and the universe isotropizes tending to a FLRW geometry. Towards the singularity, we have obtained the well-known scaling of the different scalars that show the blow-up of the curvature.
@@ Acknowledgments
This work is supported by the Basque Government Grant IT1628-22, and by the Grant PID2021-123226NB-I00 (funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”). SFU is funded by an FPU fellowship of the Spanish Ministry of Universities.

### v3

@@ 4 Explicit particular solutions
  • For cases with $\{0<\rho_{0+}$ and $0<\rho_{0-}\}$ or $\{\rho_{0+}<0$ and $\rho_{\rm threshold}<\rho_{0-}\}$, as can be seen in Figs. 8 and 8, the potential has no positive roots, and the evolution of the scale factor is given by (4.13) or (4.14), respectively. These solutions describe an indefinite expansion $a\in(0,+\infty)$, with the domain $\tau\in(-\infty,\tau_{0})$, where
    $\tau_{0}:=-\frac{1}{3P(1-\omega_{+})}\ln\left({\rm sgn}(\rho_{0+})+\frac{2\sqrt{3}P\sqrt{\rho_{0-}}}{|\rho_{0+}|\sqrt{\kappa}}\right)$
    is the time when infinite volume is reached, while the singularity is located at $\tau\to-\infty$. A particular case of this solution corresponds to the case analyzed in Ref. [3], with dust ($\omega_{+}=0$ and $\rho_{0+}=\rho_{0{\rm dust}}$), a positive cosmological constant ($\omega_{-}=-1$ and $\rho_{0-}=\Lambda/\kappa$), and stiff matter [included as $P\to(P^{2}+\frac{\kappa}{3}\rho_{0{\rm stiff}})^{1/2}$].
  • For cases with $\rho_{0-}<0$ and $P\neq 0$, the potential $V(a)$ has a single positive root (see Fig. 8), the solution is given either by (4.13) if $\rho_{0+}>0$, or by (4.14) if $\rho_{0+}<0$, with domain $\tau\in\mathbb{R}$. The range of $a$ is $a\in(0,a_{\max}]$, and thus there is a recollapse at
@@
  • For cases with $\{\rho_{0+}<0$ and $0\leq\rho_{0-}\leq\rho_{\rm threshold}\}$ the potential $V(a)$ has two positive roots (see Fig. 8),
@@
    On the one hand, it describes a finite universe with an evolution given by the functional form (4.14), with image $a\in(0,a_{-}]$, which thus presents an initial and a final singularity, as well as a recollapse at $a_{-}$, as can be seen in Fig. 12. On the other hand, it also describes a singularity-free universe, with an evolution given by the functional form (4.15) and depicted in Fig. 12, which leads to an image $a\in[a_{+},+\infty)$, and thus undergoes a bounce at $a=a_{+}$. An interesting property of this cosmology is that, even if one needs to assume an exotic fluid with $\rho_{0+}<0$ to resolve the singularity, for large volumes $a\to+\infty$ the density of such fluid decays until becoming negligible, and the other nonexotic component $\rho_{-}>0$ dominates completely the evolution. Let us also mention that, in the isotropic case $P\to 0$, the threshold energy $\rho_{\rm threshold}$ tends to infinity and, thus, the exotic $\rho_{0+}$ species is able to produce a bounce and provide a singularity-free cosmological evolution, independently of the energy contained in the $\rho_{0-}$ component.
    In the degenerate case with $\rho_{0-}=\rho_{\rm threshold}$, both roots coincide $a_{+}=a_{-}$, $V(a)$ presents a maximum there, and thus $a=a_{+}=a_{-}$ is an unstable equilibrium point. This implies that, unlike in the general case that the system reaches $a_{\pm}$ in finite time, in the degenerate case $a_{\pm}$ will be reached only in an infinite amount of time. The time evolution for this case is shown in Fig. 12.
@@ 5 Asymptotic behavior
Let us now study the asymptotic behavior, both towards the singularity and towards large volumes, of the Bianchi I model filled with several barotropic fluids. As commented in Subsec. 2.2, due to their different barotropic indices, the scaling of the species differs. For $a\to 0$, the fluid with the maximum barotropic index $\omega_{+}$ will dominate, while for $a\to+\infty$ all the species will be negligible except the one with the minimum $\omega_{-}$. Therefore, we will consider the regime where only one fluid is dominant, so we can then use the explicit solution (4.9). Since recollapsing or bouncing cosmologies with exotic matter fields may not reach the limit of either large or small volumes, for simplicity, here we will assume that in both limits the corresponding dominant species has a positive energy density $\rho_{0}$.
As commented in the previous section, we will analyze the curvature by means of the Kretschmann scalar (2.13) and the Weyl square (2.14), which are related by Eq. (2.18). Specifically, considering a universe with a single species with $\rho_{0}>0$ and $\omega\in[-1,1)$, which follows the evolution (4.9), for $P\neq 0$ the Kretschmann scalar reads
@@ 5.1 Isotropization towards large volumes
where $a_{0}:=\left[\frac{3}{4}\kappa\rho_{0}(1-\omega)^{2}\right]^{-\frac{1}{3(1-\omega)}}$, and one should take into account that the evolution (4.9) is only valid for $\omega\in[-1,1)$ and $\rho_{0}>0$, and thus also this approximation. Moreover, taking this limit in the evolution (3.6) and (3.7) of the shape parameters leads to
@@
5.2 Asymptotics towards the singularity and blowup of the curvature
@@
for $P\neq 0$ and $\omega\neq 1$. For the isotropic ($P=0$) case with $\omega\neq 1$, the evolution is given by (4.10). Hence, unless $P=0$, in this limit $a$ evolves exponentially in $\tau$, mirroring precisely the evolution of the vacuum (Kasner) solution, as given in Eq. (4.2). In order to translate this to the cosmological time $t$, we can just consider the relation obtained for the vacuum solution (4.3), which leads to the form
@@
To finish with this analysis, let us evaluate the Kretschmann scalar (5.1) in this limit. For the $P=0$ case, it scales as a power law (5.2), and therefore always diverges towards the singularity, regardless of the matter type. For $\rho_{0}>0$, $\omega\neq 1$, and $P\neq 0$ in the limit towards the singularity, $\tau\to-\infty$,
@@
As we can observe from (5.18), in terms of the cosmological time, the form of the Kretschmann scalar in this limit does not depend on the matter content. In fact, its scaling follows the same power law as in the large-volume limit (5.12). But, since the singularity is located at $t=0$, in this case the scalar will diverge.
Note that in the above expressions (5.17)–(5.19), we have only displayed the leading term, which, if $\sin(3\theta/2)=0$, will be vanishing. However, under the presence of matter, there are always subdominant diverging terms, and the Kretschmann scalar generically diverges. The vacuum scenario is the only case for which there are certain trajectories that have a constant and exactly vanishing Kretschmann scalar, which thus does not diverge towards the singularity. It is easy to see this feature from its exact form (5.3): if either $p_{+}=-P$ (and thus $\theta=0$) or $p_{+}=P/2$ (and thus $\theta=2\pi/3$ or $\theta=4\pi/3$), then $K=0$. These special trajectories correspond to the following three sets of Kasner exponents (4.6): $(p_{1}=p_{2}=0,p_{3}=1)$, $(p_{1}=p_{2}=0,p_{2}=1)$, and $(p_{2}=p_{3}=0,p_{1}=1)$, which define the only three possibilities with all the Kasner exponents being nonnegative, and it implies that none of the scale factors (4.5) expand towards the singularity.
@@ 6 Conclusions
We have obtained exact analytic solutions to describe the evolution of the Bianchi I geometry coupled to several perfect fluid species. The dynamical variables that describe the evolution of the model are the energy density of each species $\rho_{s}$, the two shape parameters $\beta_{\pm}$, which encode the anisotropy, and the average scale factor $a$, which is defined as the geometric average of the scale factors in the different spatial directions.
@@
In the chosen gauge, the evolution of the matter energy density $\rho_{s}$ and the shape parameters $\beta_{\pm}$ can be analytically obtained in general. Furthermore, the average scale factor $a$ follows the Friedmann equation (3.8), where the anisotropies are encoded in a constant term $P^{2}=(p_{+}^{2}+p_{-}^{2})$, which can be understood as the (square of the) momentum of a massless scalar field. The solution to this equation can be reduced to a quadrature, and we have described the qualitative evolution of the system by understanding the Friedmann equation as an energy equation for a particle moving on a potential [see Eq. (3.11)]. In addition, we have obtained the explicit analytic solution for a number of cases. In particular, we have derived the Kasner solution for vacuum (4.2), as well as the solution for a single fluid species (4.9). Although this solution has been studied in the literature, we have not found it anywhere given in such explicit and compact form as here. We have also considered the case with a negative energy density (4.11), which includes the model with a negative cosmological constant, and generically produces a recollapse of the universe, leading to a cosmology with an initial and a final singularity. Concerning the case with two fluid species, the solution is obviously more complicated, but, if the barotropic indices obey certain relation, the integral can be performed and the solution can be explicitly written. Such relation includes the particular cases of a Bianchi I universe either with cosmological constant and dust, or with radiation and cosmic strings. For completeness, we have obtained the explicit solution (4.13)–(4.15) for all the different signs of the energy densities. Exotic species with a negative sign of the density can produce turning points of the scale factor and bound (either from above or below) its possible values. Among the different possibilities, we find an interesting case, with an exotic and nonexotic component, which leads to a singularity-free cosmology: the universe collapses from infinite volume until a certain minimum value of the scale factor, where a bounce happens, and then it expands forever. The exotic component is dominant around the bounce, while, for large volumes, it decays and becomes negligible as compared to the nonexotic one. We consider that this, and similar models, could be worth studying as an effective description of a nonsingular cosmology.
@@
Finally, we have analyzed the asymptotic behavior of the Bianchi I universe for both large volumes and towards the singularity. In such limits, we have studied in detail the scaling of the different curvature invariants, and explicitly derive their dependence on the different parameters of the model. For large volumes, the anisotropies decay and the universe isotropizes tending to a FLRW geometry. Towards the singularity, we have obtained the well-known scaling of the different scalars that show the blowup of the curvature.
@@ Acknowledgments
This work is supported by the Basque Government Grant No. IT1628-22, and by the Grant No. PID2021-123226NB-I00 (funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”). S.F.U. is funded by an FPU fellowship of the Spanish Ministry of Universities.
Appendix A Plots for two species with $1-\omega_{-}=2(1-\omega_{+})$
Here we present certain plots for the case with two species whose barotropic indices satisfy $1-\omega_{-}=2(1-\omega_{+})$. The figures show the qualitative form of the potential $V(a)$, defined in (3.10), as well as the evolution (4.13)–(4.15) of the average scale factor $a=a(\tau)$, for different values of the densities $\rho_{01}$ and $\rho_{02}$.
Figure 5: Shape of $V(a)$ for two species whose barotropic indices satisfy $1-\omega_{2}=2(1-\omega_{1})$, with positive densities $\rho_{01},\rho_{02}>0$. In this scenario, $V(a)$ has no positive roots, and thus it corresponds to a forever expanding cosmology with $a\in(0,+\infty)$.
Figure 6: Shape of $V(a)$ for two species whose barotropic indices satisfy $1-\omega_{2}=2(1-\omega_{1})$, with negative density $\rho_{02}<0$. When $a\to 0$, the first species dominates, while for $a\to+\infty$, the second one dominates, making $V(a)$ to take positive values. Consequently, $V(a)$ has a single and positive root, $a_{\max}$. This scenario represents thus a recollapsing cosmology with $a\in(0,a_{\rm max}]$.
Figure 7: Shape of $V(a)$ for two species whose barotropic indices satisfy $1-\omega_{2}=2(1-\omega_{1})$, with negative density $\rho_{01}<0$ and positive density $\rho_{02}>\rho_{\rm threshold}$. In this case, $V(a)$ has no real roots and thus it represents a forever expanding cosmology with $a\in(0,+\infty)$.
Figure 8: Shape of $V(a)$ for two species whose barotropic indices satisfy $1-\omega_{2}=2(1-\omega_{1})$, with negative density $\rho_{01}<0$ and positive density $0\leq\rho_{02}<\rho_{\rm threshold}$. In this case, $V(a)$ has two positive roots, $a_{\pm}$, which define two regions with different cosmological evolutions: a recollapsing cosmology with $a\in(0,a_{-}]$, and a singularity-free cosmology that expands towards infinitely large volumes with $a\in[a_{+},+\infty)$.
Figure 9: Evolution (4.13) for $a=a(\tau)$, corresponding to the scenario where there are two species with barotropic indices satisfying $1-\omega_{-}=2(1-\omega_{+})$ and positive density $\rho_{0+}>0$. Each curve corresponds to a sign of $\rho_{0-}$. Depending on this sign, the cosmological evolution presents an infinite expansion (for $\rho_{0-}\geq 0$) or a recollapse (for $\rho_{0-}<0$). As can be seen, in the limit towards the singularity ($\tau\to-\infty$), both evolutions match, as it corresponds to the region where $\rho_{0+}$ dominates, and thus the value of $\rho_{0-}$ is irrelevant.
Figure 10: Evolution (4.14) for $a=a(\tau)$, corresponding to the scenario where there are two species satisfying $1-\omega_{-}=2(1-\omega_{+})$ and with negative density $\rho_{0+}<0$. Each curve corresponds to a different value of $\rho_{0-}$. Depending on the latter, there are different cosmological evolutions, presenting either an infinite expansion (for $\rho_{0-}>\rho_{\rm threshold}$) or a recollapse (for $\rho_{0-}<\rho_{\rm threshold}$). As can be observed, in the limit towards the singularity $\tau\to-\infty$, all three evolutions match, as it corresponds to the region where $\rho_{0+}$ dominates, and thus the value of $\rho_{0-}$ is irrelevant.
Figure 11: Evolution (4.15) for $a=a(\tau)$, corresponding to the scenario where there are two species whose barotropic indices satisfy $1-\omega_{-}=2(1-\omega_{+})$, with negative density $\rho_{0+}<0$ and for $0\leq\rho_{0-}<\rho_{\rm threshold}$. As can be seen, this evolution corresponds to a singularity-free cosmology, with a bounce at $a=a_{+}$.
Figure 12: Evolution of $a=a(\tau)$, corresponding to the degenerate case with two species satisfying $1-\omega_{-}=2(1-\omega_{+})$, negative density $\rho_{0+}>0$, and $\rho_{0-}=\rho_{\rm threshold}$. The singularity-free cosmological evolution, given by (4.14) and shown in purple, begins with an infinite value of $a$ and asymptotically tends to the minimum $a=a_{-}=a_{+}$. The other independent evolution, given by (4.15) and shown in blue, corresponds to an ever-expanding universe asymptotically tending to the maximum $a=a_{-}=a_{+}$.